### v1

# InterQnet: A Heterogeneous Full-Stack Approach to Co-designing Scalable Quantum Networks

Abstract
Quantum communications have progressed significantly, moving from a theoretical concept to small-scale experiments to recent metropolitan-scale demonstrations. As the technology matures, it is expected to revolutionize quantum computing in much the same way that classical networks revolutionized classical computing. Quantum communications will also enable breakthroughs in quantum sensing, metrology, and other areas. However, scalability has emerged as a major challenge, particularly in terms of the number and heterogeneity of nodes, the distances between nodes, the diversity of applications, and the scale of user demand. This paper describes InterQnet, a multidisciplinary project that advances scalable quantum communications through a comprehensive approach that improves devices, error handling, and network architecture. InterQnet has a two-pronged strategy to address scalability challenges: InterQnet-Achieve focuses on practical realizations of heterogeneous quantum networks by building and then integrating first-generation quantum repeaters with error mitigation schemes and centralized automated network control systems. The resulting system will enable quantum communications between two heterogeneous quantum platforms through a third type of platform operating as a repeater node. InterQnet-Scale focuses on a systems study of architectural choices for scalable quantum networks by developing forward-looking models of quantum network devices, advanced error correction schemes, and entanglement protocols. Here we report our current progress toward achieving our scalability goals.
Keywords InterQnet $\cdot$ heterogeneous $\cdot$ scalable $\cdot$ quantum networks $\cdot$ systems approach
1 Introduction
Quantum networks hold enormous potential for ground-breaking advances in many areas of science and technology. Distributed quantum computing [1] and sensing [2] are among the important applications envisioned for quantum networks. As with the classical Internet, further applications not yet imagined can be expected to eclipse those currently envisioned, once scalable quantum networks become a reality.
Yet despite the demonstration of entanglement distribution and teleportation of quantum states in laboratory-, campus-, and metropolitan-scale networks [3], scaling quantum networks to longer distances and to multiple users and applications remains a major challenge. The InterQnet project is advancing the state of the art in scalable quantum networks through work aimed at two interrelated overarching goals: (1) practical realization of a full-stack heterogeneous quantum network (InterQnet-Achieve) and (2) systems study of scalable quantum network architectures (InterQnet-Scale). By accomplishing these goals, InterQnet will provide a fundamental building block and foundational architectural choices for national-scale quantum networks.
Figure 1: Our systems approach to scalable quantum networks. The approach integrates devices, error correction methods, protocols, architectures, and simulation/experimentation into two co-design/integration cycles.
In InterQnet-Achieve (inner circle in Fig. 1), we are developing a repeater-based heterogeneous quantum network by co-designing (1) a set of devices including a prototype first-generation quantum repeater, microwave-optical quantum transducers, and optical quantum frequency converters; (2) related error mitigation strategies for communications; and (3) the necessary control and management protocols and interfaces, in a continuous cycle of development, validation, and optimization using physical testbeds and simulation studies.
While we use a centralized control and customized network protocols and error mitigation methods in InterQnet-Achieve, we investigate alternative quantum network architectures, sophisticated protocols, and advanced error correction methods in InterQnet-Scale (outer circle in Fig. 1) by (1) implementing them in SeQUeNCe [4]—a scalable quantum network simulator; (2) performing at-scale simulations of them by building models of future devices (in addition to the models for current devices in InterQnet-Achieve); and (3) utilizing the experimental testbeds to validate certain aspects of the simulations.
The remainder of the paper is organized as follows. Section 2 reviews the most recent experimental demonstrations of quantum networks and proposals of architectures for quantum networks, highlighting the main challenges for scalability. Section 3 describes our approach to address scalability challenges. As illustrated in Fig. 1, InterQnet involves three thrusts: Heterogeneous Devices (Section 4), Error Management (Section 5), and Network Architecture and Protocols (Section 6). These sections, along with Section 7, provide details on our progress toward the InterQnet goals. Section 8 concludes the manuscript with an outlook for future achievements.
2 Background
Researchers have demonstrated early prototypes of campus- and metropolitan-scale quantum communications [5, 6, 7, 8]. However, unavoidable signal losses make quantum repeaters essential for long-range ($\gtrsim$ 100-kilometer) communication, with repeater-assisted entanglement distribution envisioned through quantum memories [9]. Meanwhile, on the theory front, several quantum network architectures and protocol stacks have been proposed for a future quantum internet [10]. For instance, Van Meter et al. [11] proposed a layered stack approach and later developed this into a recursive architecture [12]. Wehner’s group at TU Delft has also proposed layered models for quantum networks [13, 14] that roughly map the classical TCP/IP stack. While Van Meter’s and Wehner’s work focused on bipartite entangled states, Pirker and Dür [15] extended this consideration to multipartite entangled states. Despite these advances, quantum networking still has a long way to go before it becomes a mature technology. A significant gap persists between the theory and practice of quantum networking, and one could argue that this gap may even be widening. The following subsections provide details on efforts that attempt to close this gap by building a quantum networking testbed and defining novel quantum networking architectures.
2.1 Quantum Networking Testbeds
Recent efforts to build and characterize deployed-fiber quantum network testbeds are making progress in bridging the gap between theory and practical implementation. Pompili et al. demonstrated a phase-stabilized quantum network architecture based on three entangled nodes [16] and achieved entanglement delivery [7] using the protocol stack proposed in [13]. The BARQNET testbed [17] has provided valuable insights into characterizing and compensating for channel impairments including polarization and phase noise over a 50 km fiber loop, highlighting effective integration strategies for quantum memory compatibility. The QUANT-NET testbed [18, 19] building a multinode quantum network connecting Lawrence Berkeley National Laboratory and UC Berkeley over deployed fiber, with the goal of implementing entanglement swapping and teleportation of controlled-NOT gates between distant trapped-ion processors. In the interim, QUANT-NET has deployed essential infrastructure components such as Hong–Ou–Mandel interferometry and polarization drift compensation, controlled by a two-level control framework [20].
2.2 Quantum Network Architectures
Early attempts to define an architecture and protocol stack for a quantum internet have tried to replicate the success of the classical Internet; thus, their proponents defined stacks [21, 13, 6] that map directly to the TCP/IP [22] stack. Delle Donne et al. [23] presented QNodeOS, a platform-independent operating system for end nodes of a quantum network, which enables multitasking and high-level application execution on heterogeneous quantum hardware. Besides these architectures, researchers have proposed software-defined networking-based control for quantum networks [24, 25, 26].
Departing from the state of the art, Pirker et al. [27] proposed a shift from layered architectures to a resource-centric, task-based model. In this approach, applications define high-level objectives that are fulfilled through distributed workflows—called “sagas”—that operate directly on quantum channels, classical messaging, and shared entanglement. This flexible model enables decentralized orchestration and emphasizes dynamic control, raising new questions about metrics, automation, and abstractions for scalable quantum networking. In a more radical shift, Caleffi and Cacciapuoti have proposed a novel hierarchical quantum internet architecture centered around the concept of an entanglement-defined controller [28]. Such a controller would enable scalable and efficient management of in-network operations such as the distribution and manipulation of quantum entanglement.
2.3 Open Scalability Challenges
Figure 2: Hourglass-shaped quantum networking stack. EPS is an entangled photon source.
The transition from small-scale, isolated experiments to national-scale, heterogeneous quantum networks is hindered by limitations in scalability, control, and architectural understanding. Most current efforts are restricted to homogeneous platforms and linear topologies, with little validation against real-world conditions.
Scalability in quantum networks is multidimensional. Distance remains limited by photon loss and memory decoherence. Equally important is interoperability, since practical networks must link superconducting qubits, atomic ensembles, and solid-state memories operating at different wavelengths and bandwidths. Entanglement distribution rates are constrained by probabilistic sources, multipair emissions, and imperfect detectors. Supporting multiple simultaneous users introduces new demands for orchestration and fair allocation of entangled pairs. Resilience is also critical, as accumulated noise threatens fidelity without effective error management.
Critical issues remain. Today’s networks lack well-defined abstractions for component classes, leaving it unclear how to build robust control layers across diverse technologies. Error correction and mitigation must be deployed at each node, customized to the service being provided. Architecture-level design choices are rarely guided by realistic performance and resource constraints. Interfaces that allow users and applications to access quantum network services remain poorly defined. It is also unclear how traffic engineering and quality of service can be realized when photonic states cannot be directly processed in the network.
A useful lesson from the classical Internet is not the TCP/IP stack itself but its hourglass shape. A single narrow waist connects many protocols above and below. In quantum networks, entanglement protocols can play this role, accommodating heterogeneous physical platforms underneath and diverse applications on top (see Fig. 2). Achieving such an hourglass-shaped stack will require clear abstractions for components, scalable error management, and a transversal layer for management, monitoring, and validation.
These open challenges motivate the dual goals of InterQnet-Achieve and InterQnet-Scale described in Section 3. Sections 4 through 7 then describe our approach: heterogeneous devices, error management, network architecture, and integration.
3 The InterQnet Approach
Quantum communication networks must scale along multiple dimensions, including heterogeneity of qubit platforms, distance between quantum information source and destination, number of quantum nodes in the network, rate of teleportation or entanglement distribution, and number of users. InterQnet takes a full-stack, co-design approach to enable scalable quantum networks through two overarching goals:
  • InterQnet-Achieve: Practical realization of a 3-node heterogeneous quantum network that integrates ytterbium atoms, erbium ions, and superconducting qubits, connected over the Argonne quantum network (ARQNET) [29] fiber infrastructure with quantum frequency conversion and microwave-optical transduction.
  • InterQnet-Scale: A systems study focused on architectural trade-offs critical to scalability, including centralized vs. distributed control, in-band vs. out-of-band signaling, and resource management across heterogeneous platforms.
These goals are supported by three tightly integrated technical thrusts: (1) Heterogeneous Devices, which develops and aligns physical-layer capabilities; (2) Error Management, which explores quantum error mitigation and correction strategies for communication; and (3) Network Architecture and Protocols, which designs and evaluates control and coordination across diverse topologies. Simulations are grounded in experimental measurements and conducted by using SeQUeNCe, a modular simulation framework developed by the team to support device-aware, protocol-driven architectural studies.
[…]
4 Heterogeneous Devices
A quantum repeater is a key element in a scalable quantum network. Component devices such as quantum memories and quantum frequency converters are essential to the realization of a quantum repeater. We leverage three different types of state-of-the-art qubit systems that are being developed at Argonne as quantum nodes to demonstrate a quantum network repeater (Fig. 3): trapped ytterbium (Yb) atom arrays, erbium ions (Er3+) in solids, and superconducting quantum devices. This heterogeneous architecture addresses a key scalability requirement by allowing for remote quantum entanglement and communication between different kinds of quantum information carriers in both the optical and microwave regimes.
Each of the three qubit systems can generate either photon-photon or photon-spin entanglement to serve as a quantum node for entanglement swapping and entanglement-based quantum communications. We will use the Yb atom array system as the key quantum repeater station; this will provide both long-lived quantum memories based on the atom spin states and the necessary two-qubit gates between local atoms for implementing the error mitigation schemes of Thrust 2 (see Section 5). Below, we provide detailed descriptions of the three distinct qubit platforms and clarify the role of each component device.
4.1 Ytterbium atoms
Neutral atoms trapped in optical tweezer arrays [30, 31, 32, 33] have become a compelling candidate for realizing key functions of a quantum repeater, such as high-fidelity, multiqubit entanglement distribution [34, 35]. Yb atoms are of particular interest because their long-lived, optically excited 3P0 “clock” state couples to several transitions in the telecom band (1390, 1480, and 1540 nm). These transitions enable the generation of atom-photon entangled states where the emission time (or “time bin” of a telecom photon) is entangled with the atomic state. We recently demonstrated this atom-photon entanglement using the 1390 nm 3P0 to 3D1 transition in Yb [36].
As shown in Fig. 3(a), our work will initially focus on the 3P03D1 transition at 1390 nm because of its relative simplicity [37, 36]. Recent advances in visible-to-telecom quantum frequency conversion [38] also make atom-photon entanglement mediated by visible transitions from the ground (1S0) state attractive alternatives owing to less complicated atomic state manipulation and less spontaneous emission to undesired states.
In order to achieve entanglement between two-matter qubits, the 1390 nm photons from a single atom are collected with a high numerical aperture microscope objective and sent through a long communication fiber to interfere with a photon emitted from another distant qubit system. After beam splitting and photon detection, a heralded Bell state (remote entanglement) can be generated between the atom and the distant qubit (e.g., another Yb atom, an Er ion, or a microwave qubit). With their long coherence time (>10>10\,s), Yb atoms can function as quantum memories to store the entanglement. Moreover, the atom array configuration will allow us to manipulate interactions between adjacent atoms, hence realizing quantum repeater operations. For this, two Yb atoms would be first entangled with two distant nodes, and then local operations and measurements would swap the entanglement to the distant nodes. Combined with two-qubit gates, error mitigation schemes can be implemented to improve the fidelity of entanglement swapping and quantum communication.
4.2 Erbium Ions
Erbium ions boast long spin [39, 40] and optical [41] coherence times and exhibit a natural telecom-band optical transition near the 1520–1530 nm range. These properties render them promising candidates for quantum memory applications. Our work aims to leverage recent advancements in the nanophotonic integration of Er3+ to establish a telecom-wavelength solid-state quantum memory node in ARQNET. Integration with nanophotonic cavities provides compelling advantages by significantly enhancing the single-photon emission rate via the Purcell enhancement and greatly increasing the scalability of erbium-based devices through CMOS-compatible nanofabrication.
An optical image of the Er3+ device is shown in Fig. 3(b). The device is a thin film of Er-doped titanium dioxide (TiO2) with low background spins, directly grown on top of an silicon-on-insulator (SOI) substrate. Fabricating 1D nanophotonic crystal cavities to couple with the Er3+ results in an optical photon emission Purcell enhancement factor exceeding 200, as well as a high single-photon collection efficiency ($\sim$ 40%) via fiber-to-chip coupling [42]. Similar devices of Er3+ in TiO2 thin films grown post-fabrication on SiN waveguides boast a narrow ensemble homogeneous linewidth of 5 kHz and spectral diffusion of 27 kHz over 4 ms, well within the 1 MHz microwave-optical (M-O) transducer bandwidth [43].
To entangle Er3+ with other qubits, we will use our cavity devices to selectively engineer a cycling transition [44, 45]. The initialization and preparation of a superposition spin state of the erbium electron can be accomplished by applying an excited state microwave (MWe) $\pi/2$-pulse followed by the emission of a telecom photon. Subsequently, a spin-selective optical $\pi/2$-pulse can be applied to rotate the spin state, thereby generating an entangled spin-photon state.
4.3 Superconducting quantum devices
A scalable quantum communication network should be able to not only transmit quantum information carried by optical photons but also interface with solid-state qubits working at the microwave frequencies [10, 46]. At the microwave quantum node, we will leverage GHz qubit devices based on superconducting transmons and/or single electron-on-solid neon (eNe) qubits [47] to create and receive microwave quantum communication signals and integrated quantum transducers to generate M-O photon entanglement [48, 49], hence achieving remote entanglement between a microwave qubit and a distant Yb atom or Er ion.
[…]
4.4 Optical Quantum Frequency Converters
Quantum frequency converters (QFCs) provide the important capability to convert an input mode of light from one frequency to another while preserving its quantum state. A QFC was proposed and demonstrated by Kumar and coworkers [55, 56] in the 1990s and has since been recognized as a key component for emerging quantum networks. In our heterogeneous network, different quantum nodes emit photons at different wavelengths (1390, 1520, 1550 nm for Yb, Er3+, POM, respectively). Thus, we will develop high-efficiency QFC devices for converting single photons emitted from different qubit systems to a matched wavelength of 745.7 nm, enabling two-photon interference and Bell state measurements to herald remote entanglement.
We have designed QFC schemes in periodically poled lithium niobate waveguides with a strong $\chi^{(2)}$ nonlinearity (see Fig. 3(d)). We focus on bulk-optic designs to keep coupling losses low and will use cavity enhancement to match the interaction bandwidth to those of the incident photons. By using pump wavelengths in the telecom bands, we will be able to convert any pair of wavelengths specified above to the common wavelength of 745.7 nm through sum frequency generation. Besides bandwidth, conversion efficiency, and added noise, another important co-design parameter is the temporal shape of the single photons, which needs to be matched between different quantum emitters to remove photon distinguishability. We have previously shown that such wavefunction engineering can be done by controlling the temporal shape of the pump beam [57, 58] during quantum frequency conversion. The capability of temporal mode shaping will play an important role in our heterogeneous network, allowing us to maximize visibility of two-photon interference and hence optimize Bell state measurements.
4.5 Summary of Achievements and Open Challenges
For Yb atoms, we have demonstrated high-fidelity atom-photon entanglement directly in the telecom band—a key building block for remote entanglement. The main limitation today is photon collection efficiency. In the near term we are improving rates with parallel multiplexing; in the longer term we will couple atoms to optical cavities to enhance collection. In Er-doped TiO2, we have achieved a $1000\times$ narrowing of homogeneous linewidths at sub-Kelvin temperatures. This is critical for isolating single ions and achieving long coherence. Next steps include fabricating high-Q cavities and verifying performance with single emitters for telecom-compatible quantum memory. For superconducting devices, we are nearing completion of two key components: SNAIL-based readout resonators and piezo-optomechanical transducers with phononic crystal engineering. Upcoming work will focus on tuning nonlinearity and boosting mechanical Q to improve conversion efficiency. In frequency conversion, we have recently completed classical characterization of the devices, and we are now investigating noise sources—especially Raman noise—to support quantum-compatible conversion. Altogether, these efforts represent diverse but converging progress toward scalable, hardware-aware entanglement distribution across heterogeneous platforms.
We are creating models for the above-mentioned key quantum devices in SeQUeNCe and applying numerical optimization methods to search for critical device parameters to maximize the fidelity of simulated networks. We successfully modeled the quantum transducer in SeQUeNCe [59]; the other key devices are works in progress. Via simulations, we will not only better understand the behaviors of current experiments but also obtain insights for future device designs and improvements. We will conduct “what-if” studies that will help capability development in future devices. For example, what if the dark count in the detectors is reduced by a factor of 10, or atomic operations can run an order of magnitude faster? Furthermore, we will develop specialized numerical optimization methods to optimize operation conditions for high-fidelity quantum communications in future large-scale quantum networks.
5 Error Management
Different from qubit-based quantum computing, optical quantum communication relies on bosonic modes to generate entangled states for long-distance transmission, without the need for universal gate sets. Thus, we require methods that can tolerate photon loss, imperfect entanglement generation, and heterogeneous device noise, while remaining practical for near-term testbeds and scalable to future large networks. InterQnet adopts a layered approach to error management: (1) lightweight error detection via Pauli Check Sandwiching (PCS) and postselection, (2) resource-aware strategies that leverage gauge freedom to simplify correction, (3) entanglement distillation protocols that convert multiple noisy Bell pairs into fewer high-fidelity ones, and (4) advanced quantum LDPC (qLDPC) and bosonic codes that enable scalable, fault-tolerant entanglement purification. Together, these techniques form a hierarchy that links the device-level realities of Section 4 to the architectural protocols of Section 6, and they map naturally onto the two project goals: enabling practical demonstrations in InterQnet-Achieve and guiding at-scale design choices in InterQnet-Scale.
5.1 Pauli Check Sandwiching
Figure 4: (a) Pauli checks forming distance-2 code. Red checks can be applied recursively. (b) Noisy simulation for different levels of recursion: rec0 represents no red gates; rec1 represents adding ancillas $a_{3}$ and $a_{4}$ and the supported red gates to the rec0 circuit; rec2 corresponds to adding ancillas $a_{5}$ and $a_{6}$ and the supported red gates to the rec1 circuit.
In previous work we investigated quantum error detection based on Pauli check sandwiching (PCS) [60], which shares many properties with quantum codes. In PCS the unitary circuit $U$ to be error mitigated is required to respect the symmetry $P_{L}UP_{R}=U$, where $P_{L}$ and $P_{R}$ are Pauli strings of the Pauli group. The control of Pauli operators is on ancillary qubits, and error detection is performed by measuring these ancillary qubits and postselecting on all zero outcomes. PCS provides protection on the targeted qubits and generally incurs less resource overhead than standard quantum error correction and detection codes do. From the point of view of quantum codes, PCS without checks on the ancillas is generally a distance 1 code (i.e., cannot detect arbitrary single-qubit errors), but we can detect arbitrary single-qubit errors on the targeted qubits by using the generators of the Pauli group as checks [60]. Additionally, there is a direct correspondence between quantum codes and entanglement purification [61, 62]. Thus, the performance of PCS in quantum networks is an important area of investigation.
The PCS scheme [60] for quantum networks performs PCS in a distributed fashion. In the simplest quantum repeater network, the PCS scheme can be executed as follows. Two nodes, $A$ and $B$, want to establish entanglement via a repeater $C$. AA and $B$ prepare GHZ-like states on their ends and send over two qubits each (one computation register qubit and one ancilla) to $C$ via an optical fiber link. These form the left end of the PCS scheme. The repeater $C$ then applies the right end of the PCS scheme along with entanglement-heralding measurements and then does postselection by performing measurements on the ancillary qubits. This can be enhanced by adding additional checks and ancillas. The novel contributions of our current work are as follows [63]. Assuming depolarizing errors, we derived the analytical equations for the postselected fidelity and postselection rate as a function of the noise for PCS $X$ checks and PCS X&ZX\&Z checks. For PCS $X$ checks we have
$c=\frac{1}{9}(1+2F)^{2}$ (1)
and
$F^{\prime}=\frac{9F^{2}}{(1+2F)^{2}},$ (2)
and for PCS X&ZX\&Z we have
$c=\frac{1}{324}(3+6F-\sqrt{12F-3}+4F\sqrt{12F-3})^{2}$ (3)
and
$F^{\prime}=\frac{1+52F^{2}-\sqrt{12F-3}-2F(4+\sqrt{12F-3})}{(\sqrt{12F-3}-1-8F)^{2}},$ (4)
where $c$ is the postselection probability, $F^{\prime}$ is the postselected Bell state fidelity, and $F$ is the initial fidelity (i.e., fidelity of the Bell state without PCS). Next, we simulated our scheme assuming depolarization noise on the optical fiber channel and witnessed fidelity gains. The fidelity gains also appeared in simulations with noisy gates. PCS can be enhanced by adding $Z$ checks on the ancilla qubits. We can apply this process recursively by also checking the new ancillas with $Z$ checks. We discovered that this process generates a family of distance 2 codes. The structure of this family is described in Fig. 4(a), where the red gates can be applied recursively. Since the code distance is fixed, only the first recursion of ancilla checks provides significant improvements in fidelity. We also simulated this method with noisy gates and attained fidelity improvements as shown in Fig. 4(b).
5.2 Quantum Resource Correction
In certain cases we may be interested only in correcting certain properties of a quantum state. Intuitively, this leads to less resource overhead than standard quantum error correction codes (QECCs), which aim to correct the logical state exactly. Thus, we investigated correcting the amount of a quantum resource in [64]. In quantum networks, entanglement is the most commonly used resource. We derived necessary and sufficient conditions for logical coherence correction. While the conditions are equivalent to the QECC conditions, recovery is simplified. Resource-preserving operations can result in gauge freedom in the code space. For instance, for the $l_{1}$ norm measure of coherence [65] and incoherent basis in the standard logical basis, logical Pauli operations are gauge transformations. In contrast, in standard QECC, unwanted logical Pauli operations are logical errors. Thus, the optimization in decoders for stabilizer codes such as the surface code is simpler for resource correction than in QECC. This is an important result since the decoding problem is in the class of NP-complete or #P-complete problems, depending on whether degenerate decoding is used [66]. Importantly, fast decoders can have the same accuracy as generally more accurate decoders when applied to resource correction. These results extend to logical entanglement and other resources that have resource-preserving operations that have a similar structure. We also investigated extending the logical incoherent basis over the physical space and identified channels that preserve the coherence of any state in the code space. Consequently, stabilizer measurements can be combined in many situations.
5.3 Entanglement Distillation and Error Detection/Correction
We optimize the entanglement purification protocol to reduce resource overhead while enhancing overall performance. In principle, entanglement distillation can asymptotically extract multiple copies of perfect Bell pairs, limited only by the distillable entanglement [67, 68]. In practice, with finite noisy entangled inputs, most purification protocols reduce noise and concentrate entanglement in the output but do not yield perfect Bell pairs. For instance, certain protocols require four noisy states as input to suppress depolarizing errors in the output state effectively. From an engineering perspective, these improved entangled states from finite-size protocols are valuable for constructing nontrivial applications.
By exploiting the correspondence between quantum error correction/detection and entanglement distillation [69], we can devise a wide array of efficient entanglement purification protocols, assisted by one-way or two-way classical communication, based on quantum error-correcting codes. For example, Fig. 5(a) depicts a quantum circuit based on the $[[4,2,2]]$ code, also known to be equivalent to the Leung–Shor protocol [70]. Upon success, it transforms four noisy inputs into two purified entangled states with higher fidelity.
To ensure practicality, we analyze the circuit implementations of entanglement purification on popular platforms. Real physical systems often impose additional constraints that hinder the efficient execution of the desired protocols. In Rydberg atom array systems, leveraging their unique features, we demonstrate the robustness and universality of efficient entanglement purification protocol implementations [71].
This advancement enhances our understanding of reliably establishing entanglement over long distances. It also facilitates the development of advanced quantum repeater architectures across various generations [72, 73, 74]. We are conducting simulations and evaluating novel protocols using SeQUeNCe, the Stim [75] Python package, and QETLAB [76] to understand the scalability trade-off of deploying these protocols in quantum networks.
5.4 qLDPC-Based Entanglement Distillation
Figure 5: (a) Quantum error detection for novel entanglement distillation protocol that can distill two Bell pairs from four imperfect ones. (b) Quantum error correction with concatenated bosonic GKP code and qLDPC code. (c) Efficient and robust purification of logical Bell pairs with qLDPC codes.
We are designing robust bosonic quantum error-correcting codes specifically tailored for quantum communications, which rely on optical wave packets for communication over long distances. We aim to enhance the performance of quantum error correction by optimizing both the choice of the logical subspace and the recovery operation. Previous investigations have shown that the single-mode Gottesman–Kitaev–Preskill (GKP) code achieves optimal performance for encoding with a single bosonic mode, effectively mitigating dominant excitation loss errors in quantum communication [77]. Furthermore, a recent experiment has demonstrated the above-break-even performance of the GKP code [78], which is highly encouraging. However, open questions remain regarding the exact ultimate performance limit and the achievement of optimized recovery, even for single-mode GKP encoding: Can we restore the encoded quantum information with fidelity above 99.99% even with a 10% loss rate? Moreover, extending this approach practically to multiple bosonic modes to approach the quantum channel capacity remains unclear. To address these challenges, our approach involves two key steps. (1) We will use the transpose-recovery-channel method (known as Petz recovery map) to explicitly construct a recovery operation that is close to optimal for the GKP code. This will not only provide an asymptotic performance limit for single-mode GKP encoding but also improve the design of the experimental GKP-QEC recovery operation. (2) We will investigate the concatenation of bosonic quantum codes with the recent development of good quantum low-density parity-check codes [79, 80], which exhibit favorable scaling in both coding rate and distance rate (see Fig. 5(b)). This will enable the development of multimode encoding schemes that approach the quantum channel capacity for future quantum networks [81].
Recent progress in entanglement purification involves the use of various quantum low-density parity-check (qLDPC) codes [82]. Fig. 5(c) illustrates the purification of Bell states at the logical level. The low density of stabilizer checks mitigates undesirable error propagation. In simulations of noisy quantum circuits, as the system size increases, we observe improved purified entanglement resources. Its scalability and low overhead of encoding and decoding render it an ideal candidate subroutine for future quantum networks.
5.5 Summary of Achievements and Open Challenges
We have investigated several complementary approaches to error management. Pauli Check Sandwiching (PCS) improves fidelity through Pauli operator checks and postselection, though implementing the necessary entangling gates remains challenging. Quantum Resource Correction (QRC) exploits gauge freedom to construct resource-preserving recovery maps that simplify decoding relative to conventional QECC. Entanglement distillation protocols, designed for dual-species atom arrays and compiled into circuits with only global control, enable purification on hardware with limited addressability, though coupling and measurement performance must improve. At a more advanced level, quantum LDPC codes leverage sparse stabilizers to achieve fault-tolerant purification with high thresholds, while multimode bosonic codes such as GKP, combined with Petz recovery, can approach channel capacity [83, 84]; concatenating these methods provides a scalable path toward efficient entanglement distribution.
Ongoing work is expanding these methods to more general noise models and developing collective bosonic encodings, offering a range of scalable, hardware-aware solutions for robust entanglement distribution. A key near-term goal is to demonstrate improved Bell violations on ARQNET using PCS. Violating a Bell inequality marks a transition away from local hidden-variable models and provides a clear experimental milestone for InterQnet-Achieve.
Together, these approaches provide a layered strategy. Lightweight PCS, QRC, and small-code distillation enable practical demonstrations in InterQnet-Achieve, while advanced qLDPC and bosonic codes inform scalability analyses in InterQnet-Scale.
6 Network Architecture and Protocols
Figure 6: Architecture of a scalable quantum network using a plane abstraction.
Fig. 6 illustrates a scalable quantum network organized using a three-plane abstraction. The infrastructure plane at the bottom represents all physical devices and fiber connections. The logical plane in the middle captures the repeater-chain abstraction once a network orchestrator has identified a suitable path between end nodes (e.g., A and B in dark green). The control and service plane at the top encompasses orchestration, routing, and user-facing services accessed through standard interfaces. We adopt this layered abstraction to structure our work on network architecture and protocols, enabling systematic study of scalability across distance, node and user counts, heterogeneity, and entanglement distribution rates, and providing a framework for integrating device models from Section 4 and error management methods from Section 5 into protocol-level evaluations.
6.1 Faithfully Simulating the Infrastructure Plane of Quantum Networks
Given the heterogeneous nature of InterQnet, SeQUeNCe has been designed to work with multiple architectural paradigms of the infrastructure plane. For this iteration of the work, we report simulation modes to tackle two important aspects of quantum communications: optical and atomic communications.
6.1.1 Simulating Optical Quantum Communications
The primary challenge in simulating fiber-optic quantum communication lies in the large Hilbert space required to model the many optical modes involved. Simulating open quantum systems with density matrices for such large systems is computationally expensive. Even with their sparsity, the exponentially increasing memory demands of sparse density matrices quickly make simulations intractable for systems with high local Hilbert space dimensions and multiple modes. To address this situation, we employed tensor networks to efficiently represent large state vectors as matrix product states [85].
Fig. 7 compares the memory requirements for simulating a polarization-based entanglement swapping experiment [86] using tensor networks versus sparse density matrices, showing a substantial reduction in memory usage as the Hilbert space dimension increases.
Figure 7: Simulation memory vs. local Hilbert space dimension. The simulation was performed over multiple cut-off values of tensor network decomposition errors for two different mean photon numbers (MPN).
We employ quantum trajectory theory [87]—which simulates open quantum system dynamics—to model the evolution of system statistics throughout the experiment. Kraus operators corresponding to the relevant noise models (e.g., amplitude damping, amplification, and depolarization) are applied to the matrix product states as they evolve through the system. These Kraus operators are applied probabilistically, with probabilities determined by the unnormalized weights of each trajectory [88]. Computing these probabilities for every iteration of a Monte Carlo experiment is computationally expensive and inefficient. Therefore, we developed a novel trajectory caching scheme. Since most quantum channels have certain trajectories that occur far more frequently than others, the required calculations for these common trajectories are performed only once, after which the cached results are reused for the majority of Monte Carlo samples.
6.1.2 Simulating Physical Qubits and Their Emission
Another challenge of faithful quantum network simulation lies in simulating entanglement generation protocols using a realistic hardware model of physical qubits. The fidelity and success rate of a protocol for generating entanglement depend on the hardware specifications, which consist of two distant qubits connected via an optical fiber link. A realistic simulation is essential for accurately predicting results that are comparable to experimental outcomes. We constructed a model in the Lindblad form to account for the internal dynamics of physical qubits and photon emission into optical fibers, as well as decoherence and imperfections. We are developing a module that enables SeQUeNCe users to precompute the outcomes of entanglement generation protocols for use in the network simulation. The module aims to provide sufficient flexibility for various user-defined physical qubits and operational sequences of protocols. We are also exploring possible ways to integrate this physical qubit simulation with the fiber-optical simulation previously described.
6.2 Entanglement witnessing in the logical plane of quantum networks
Resilience is typically defined as the ability to survive either a malicious or unintentional network failure. To fully ensure resilience in the logical plane of scalable quantum networks, one must ascertain that the distributed entangled states are actually shared between the desired users. For instance, noise could deteriorate the entanglement held in quantum memories of a repeater before a swap operation is scheduled. Or more dramatically, two end nodes may have established entanglement with a malicious repeater [89]. In both cases, we want the entanglement distribution schemes to be resilient and amenable to verification. Therefore, we consider the problem of efficiently witnessing or certifying two or more sources of entanglement in a network by performing joint measurements instead of just single-qubit measurements. By using joint measurement at network nodes, it becomes possible to simultaneously certify the distribution of entanglement from multiple sources using fewer shots than certifying the sources independently.
Our solution is to use tools from quantum self-testing and semi-device-independent quantum information processing to develop entanglement witnessing tests. Similar to how the famous CHSH inequality can certify that entanglement has been distributed to two distant locations, our work will develop similar inequalities or tests that can certify the entanglement of two or more sources on a network. Prior work has already shown the utility of this approach in special scenarios [90], but the general methodology applied to simultaneous entanglement witnessing is relatively unexplored. We anticipate making new progress by invoking prior results in the study of local operations and classical communications to help develop entanglement witnesses. One key initial accomplishment is the development of a certification test that can simultaneously witness two noisy entangled states that are distributed in a three-node line network, known as the bilocal network. While this result is not significantly different from prior known results [91], it is a starting point for this research. The biggest challenge is that the optimization problems we encounter are non-convex, and therefore standard numerical techniques are not efficient. We are exploring certain relaxations that can make the numerical optimizations more tractable.
6.3 Architectural Approaches to the Control and Service Plane of Scalable Quantum Networks
In prior work [92] we proposed a theoretical framework that enables the continuous generation of entangled pairs between neighboring nodes. This approach can adapt to incoming client requests to reduce the time to serve future requests. In [92], however, the theoretical framework is evaluated only through small-scale, ad hoc, and naive time-slotted simulations. The challenge is to implement the theoretical framework into a comprehensive quantum network protocol and simulate the protocol in a discrete-event quantum network simulator at a large scale, including realistic physical parameters.
The challenges are addressed in [93], where we designed and implemented the Adaptive, Continuous entanglement generation Protocol (ACP). ACP reduces the time to serve a client request by letting a node attempt to generate entangled pairs with neighboring nodes in the background continuously, regardless of incoming requests. The selection of neighbor nodes adapts to the request traffic by a feedback control algorithm. Additionally, entanglement distillation is used to enhance the fidelity of entangled pairs. We extended the SeQUeNCe quantum network simulator to support the simulation of ACP. In particular, we added a single-heralded entanglement generation protocol that accounts for pregenerated entangled pairs.
Figure 8: Simulation results for the 200-node autonomous system topology network.
We showed promising results when we evaluated the ACP in the SeQUeNCe simulator under various scales (2-node, 20-node, and 200-node networks) and realistic physical parameters (Table 1 in [93]). For instance, Fig. 8 shows the performance of ACP in the 200-node autonomous system topology network, comparing with on-demand-only (ODO) and uniform-continuous protocol (UCP). Fig. 8(a) shows the effectiveness of the continuous entanglement generation by UCP having a slightly lower time to serve than ODO. Fig. 8(a) also shows the effectiveness of the adaptiveness of ACP by having a 50+% lower time to serve than UCP. Adapting to incoming requests is a process that takes time, and explaining ACP’s red line gradually decreases. The sudden change in the request traffic pattern results in a spike for ACP’s time to serve (red line). Then, ACP will gradually adapt to the new request traffic pattern. Fig. 8(b) shows that ACP enhanced with entanglement distillation increases the fidelity of the end-to-end entangled pairs by as much as 0.05 compared with UCP and 0.04 compared with ODO, both without entanglement distillation.
6.4 Summary of Accomplishments and Open Challenges
The three-plane abstraction provides a unifying framework for reasoning about quantum network architectures. The infrastructure plane captures heterogeneous devices and links, the logical plane describes the generation, purification, and verification of entanglement, and the control/service plane coordinates these resources to support user applications. Our architectural analysis highlights that scalability depends not only on distance but also on node and user counts, heterogeneity, and entanglement distribution rates. These studies are designed to incorporate the error management methods of Section 5 as implementations mature, enabling evaluation of purification, distillation, and error correction in the context of realistic network protocols.
We have made significant progress in creating faithful and computationally efficient simulations of the infrastructure plane of quantum networks. With these tools, we discovered that visibility of entanglement swapping varies depending on measurement basis and source brightness, an effect arising from multipair emissions and realistic threshold detection. Our adaptive continuous entanglement generation protocol significantly improves the performance of scalable quantum networks, reducing the time to serve by up to 94% and improving fidelity by 0.05 in large-scale simulations. We are now scaling and optimizing this for hybrid links, larger networks, and centralized vs. decentralized control paradigms. Furthermore, we have made progress on SeQUeNCe integration, enhancing physical layer realism (see Section 4) and implementing protocol-level optimizations. Future work includes integrating devices and error correction models and comparing centralized vs. distributed control schemes in the control and service plane of quantum networks.
Collectively, these efforts build a scalable, realistic foundation for deploying and evaluating next-generation quantum networks. They provide the architectural context for InterQnet-Achieve demonstrations in ARQNET and for InterQnet-Scale simulations that explore qLDPC-based purification and large-scale orchestration. Section 7 builds on these results by describing how architectural protocols are realized in practice through integrated orchestration and control.
7 Integration
Integration efforts on ARQNET bring the architectural concepts of Section 6 into practice, focusing on the coexistence of quantum and classical signals, stabilization of quantum channels under realistic fiber conditions, and orchestration of heterogeneous devices. These activities are essential to InterQnet-Achieve, where a three-node heterogeneous network will demonstrate entanglement distribution across superconducting, erbium, and ytterbium platforms, and they also inform InterQnet-Scale by identifying key constraints and performance trade-offs for realistic architectures.
7.1 Quantum-Classical Coexistence over Deployed Fiber
We prototyped an on-demand entanglement distribution service over ARQNET, addressing the challenge of spontaneous Raman scattering (SpRS) induced by coexisting classical signals [29]. We developed an empirical model that quantitatively describes the relationship between the coincidences-to-accidentals ratio (CAR) and pump power of entangled photon sources in the presence of classical traffic:
$CAR=\frac{\alpha_{s}\cdot\alpha_{i}\cdot\mu_{c}}{((\mu_{c}+\mu_{sn})\cdot\alpha_{s}+d_{s})((\mu_{c}+\mu_{in})\cdot\alpha_{i}+d_{i})}+1,$ (5)
where $\alpha_{x}$ denotes transmittance, $\mu_{c}$ represents correlated photon pairs per pulse, and $\mu_{xn}$ represents noise photons from SpRS. This framework enables operators (or a future network orchestrator) to optimize quantum device settings dynamically based on classical traffic levels and achieve a maximum CAR under the defined constraints.
Figure 9: Dependence of the coincidences-to-accidentals ratio (CAR) on average power of the pump pulses for varying amounts of SpRS.
We established a proportional relationship between our theoretical model and the actual input power of an entangled photon source, which we subsequently utilized to execute experimental demonstrations of the on-demand entanglement service. Experimental validation was conducted on ARQNET utilizing deployed fiber connections between buildings 360 and 440, covering 4.38 kilometers (round trip). The experimental setup integrates classical traffic generation using 10GBASE-LR transceivers at 1330 nm with a commercial entangled photon source (NuCrypt EPS-1000-W) operating at 1550.10 nm. Classical and quantum signals are multiplexed using wavelength division multiplexers and routed through a software-controlled optical cross-connect, enabling both co-propagation and counter-propagation scenarios (only co-propagation is presented).
Entanglement quality was validated through two-photon interference measurements using a 1 ns coincidence window, achieving coincidence rates of 155–120 counts per second. Fringe visibilities calculated as $V=(N_{max}-N_{min})/(N_{max}+N_{min})$ ranged up to 68% and remained within a similar range with or without classical light being present in the system. Our simulations and preliminary tests at ARQNET demonstrate the model’s capability to optimize quantum network performance by guiding configuration parameters such as EPS pump power, detector bias, and classical signal intensity. However, increasing the number of automated variables may introduce additional forecasting complexities for the model.
Figure 10: (a) Schematic of the experimental setup showing the integration of classical and quantum traffic over deployed fiber infrastructure with multiplexing and filtering components. (b) Two-photon interference fringes showing maintained entanglement quality with visibilities of 59–68% in the presence of coexisting classical traffic.
These results validate the feasibility of automated entanglement distribution services where users can request specific CAR values and the system dynamically adjusts pump power settings based on the predictive model. The successful coexistence of quantum and classical signals over deployed fiber infrastructure establishes a foundation for scalable quantum networking architectures integrated with existing telecommunications infrastructure. We also demonstrated our quantum network’s capability to achieve Hong–Ou–Mandel interference between photons generated at two distinct sources and transmitted through deployed fiber, alongside a coexisting clock signal generated by transceivers running at 2 Gbps [94]. We verified this interference between weak coherent-state photons and photons from a heralded single-photon source located remotely. A maximum dip visibility of 0.58 $\pm$ 0.04 was achieved. The sources were connected via 4.3 km of deployed fiber and had a coexisting clock signal in the same fiber to synchronize them. This represents a first step toward achieving teleportation over our network and demonstrates that it maintains sufficient timing and polarization stability required for scalable quantum networking.
7.2 Polarization Stabilization Using Machine Learning
Random variations in the birefringent profile of optical fibers—due to both static manufacturing defects and dynamic environmental processes—cause rotations to the state of polarization (SOP) of the electric field. While orthogonal states will remain orthogonal, in general, the polarization state at the output end of a fiber channel will not match the one sent at the input channel. Thus, in order to properly interfere quantum information photons encoded in polarization, two nodes must align their polarization bases so that they match.
We are currently studying the use of machine learning to find correlations between quantum information signals and classical control signals separated in wavelength. In the past, limitations caused by polarization mode dispersion (PMD) and the corresponding differential group delay, $\tau$), have required that control signals be placed within $\approx 2$ nm of their quantum counterparts [95]. We used an attention-based machine learning model to look for higher-order correlations between distantly separated signals. We generate training data for the model with a polarization scrambler, consisting of a motorized polarization controller to vary the SOP of the two signals followed by a 25 km spool of SMF-28 fiber to consider the effects of PMD (see Fig. 11(a)). After demultiplexing the two signals using a wavelength division multiplexer (WDM), we send each signal to a Stokes polarimeter, which reports each SOP as three Stokes parameters, which can be considered Cartesian coordinates on the Poincaré sphere. Our model directly predicts the SOP of the data signal based on the SOP of the control signal. Our most recent results [96] were able to predict the SOP of a signal separated from the control by 5 nm with an average root mean square error of 0.072 in individual Stokes parameters, as demonstrated in Fig. 11(b).
Figure 11: (a) Experimental setup to collect SOP polarization data for machine learning. The polarization controller acts as a scrambler, and the fiber spool is used to add the effect of polarization mode dispersion. (b) Main: Measured S1 parameter and our model predictions. The model faithfully tracks the state of polarization with a root mean square error of 0.072. Inset: Measured SOP states on the Poincaré sphere, demonstrating that we sample all possible SOPs.
More flexibility is granted for allocating polarization control signals by allowing greater separation between the control and quantum signals. In a perfect scenario, any free WDM channel in the network can be used as a polarization control channel, so long as the machine learning model can accommodate the degree of separation between the two signals. Moreover, if a control channel needs to be reallocated, it can be moved to any other open channel without degrading the quality of polarization tracking. Furthermore, by training the model on a number of different wavelength separations, one might be able to stabilize multiple quantum channels with a single control signal, reducing the total number of control channels required in a network.
7.3 Quantum Network Orchestrator Prototype
To coordinate heterogeneous quantum devices and protocols on ARQNET, we developed a modular orchestrator that instantiates the three-plane abstraction introduced in Section 6. The orchestrator adopts a microservice architecture where each physical device such as entangled photon sources, polarization analyzers, or time taggers is encapsulated as a gRPC-based software agent. This design enables the orchestrator to issue high-level commands while device-specific control logic is handled locally, allowing protocols such as two-photon interference or quantum state tomography to be executed automatically across distributed nodes. Fig. 12 illustrates the orchestrator architecture, showing the gRPC-based interaction between device agents and the central control service.
A central challenge in distributed photonic networks is synchronization. Our prototype integrates a lightweight hybrid clock distribution scheme that simultaneously transmits 10 MHz and 1 PPS signals via radio-over-fiber, combined with NTP-based alignment. This approach provides picosecond-level timing coordination between remote Swabian time taggers without the complexity of GPS or White Rabbit systems, ensuring that coincidence detection and entanglement verification can be performed accurately across kilometers of fiber.
The orchestrator also incorporates polarization drift compensation as a gRPC service, enabling autonomous calibration and stabilization routines that replace manual realignment. To our knowledge, this is among the first orchestrators to implement a plane-based abstraction in a deployed quantum network testbed.
Figure 12: InterQnet orchestrator prototype following the plane abstraction.
7.4 Summary of Accomplishments and Open Challenges
Integration efforts on ARQNET have demonstrated feasibility across multiple layers. Coexistence studies showed that entanglement can be maintained alongside classical traffic, supported by an empirical model of Raman scattering and validated experimentally over deployed fiber. Polarization stabilization experiments using machine learning demonstrated that quantum channels can be predicted from classical control channels, with the potential to reduce the number of wavelengths needed for robust operation under dynamic conditions. A prototype quantum network orchestrator was also developed, implementing the three-plane abstraction with serial drivers for device control and gRPC-based coordination.
Challenges remain in extending coexistence studies to more traffic patterns and longer spans, demonstrating polarization stability over extended timescales, and scaling the orchestrator to manage heterogeneous entanglement across superconducting, erbium, and ytterbium platforms. These integration efforts complement the error management strategies of Section 5 and the architectural analysis of Section 6, advancing InterQnet toward its goals of heterogeneous demonstrations (Achieve) and guiding scalable design principles (Scale).
8 Conclusion
The path from small-scale demonstrations to national-scale, heterogeneous quantum networks remains blocked by challenges of scalability, control, and architectural clarity. Most existing efforts focus on homogeneous systems or simulations disconnected from experimental feedback. InterQnet distinguishes itself by: (i) integrating heterogeneous platforms, including atoms, ions, and superconducting qubits, in the first demonstrations of their kind; (ii) performing architectural simulations in SeQUeNCe that are parameterized by realistic device and network constraints, informed directly by experiments; (iii) pursuing full-stack co-design, where protocols, control logic, and device models are developed in tandem to enable end-to-end system exploration; and (iv) evaluating architectures under realistic conditions, exploring trade-offs in service time, fidelity, and resilience within fiber-deployed topologies rather than abstract models.
Through this combination of experimental integration and systems-level study, InterQnet has advanced toward building a full-stack heterogeneous quantum network (InterQnet-Achieve) and guiding the design of scalable architectures (InterQnet-Scale). This work establishes a foundation for future demonstrations such as teleportation and Bell inequality violations on ARQNET, as well as scaling qLDPC-based purification to larger networks. Together, these results move the field closer to the realization of robust, heterogeneous quantum networking at national scale.